OpenStudy (anonymous):

each interior angle of a regular polygon is 90 less than five times the measure of each exterior angle. how many sides does the polygon have

OpenStudy (calculusxy):

let the value of each interior angle be x then the value of each exterior angle will be 180-x now it is given that each interior angle is 90 less than five times the measure of each exterior angle. so x=5(180-x)-90 => x=900-5x-90 => 6x=900-90 => x=150-15 => x=135 now the value of the interior angle of a regular polygon of n sides is given by 180(n-2)/n degrees so here 180(n-2)/n = 135 => 180n-360=135n => 45n=360 so n=8 The no. of sides of the polygon = 8 so it is an octagon Source: algebra.com

OpenStudy (anonymous):

how did it become 150?

OpenStudy (calculusxy):

Where did you get 150?

OpenStudy (anonymous):

6x=900-90 x=150-15 ?

OpenStudy (anonymous):

it's supposed to be 810?